# The effect of accretion on scalar superradiant instability

### v1

@@ I Introduction
Superradiance allows ultra-light bosons to extract energy and angular momentum from rotating black holes (BHs). This mechanism exists when the condition $\omega<m\Omega_{\mathrm{H}}$ is satisfied, where $\omega$ is the frequency of the boson field, $m$ is the magnetic number, and $\Omega_{\mathrm{H}}$ is the angular velocity of the event horizon. This process leads to a significant accumulation of bosons in bound states surrounding a Kerr BH, forming a BH-condensate system analogous to a hydrogen atom. For a comprehensive review on superradiance, we refer the readers to Ref. Brito:2015oca . In this work, we study the superradiance of ultra-light scalars, such as the QCD axion Weinberg:1977ma ; Wilczek:1977pj , and axion-like particles suggested by string theory Arvanitaki:2009fg .
@@
Many observables are proposed to observe such BH-condensate systems Chen:2019fsq ; Chen:2021lvo ; Blas:2020nbs . In this article, we focus on the gravitational effects of the scalars, independent of how the scalar field couples to the Standard Model particles. Both direct and indirect detection methods have been proposed in the literature. One could directly observe the gravitational wave (GW) signals from the BH-condensate system, including quasi-monochromatic GWs Arvanitaki:2010sy ; Yoshino:2013ofa ; Yoshino:2014wwa ; Arvanitaki:2014wva ; Arvanitaki:2016qwi ; Brito:2017wnc ; Brito:2017zvb ; Baryakhtar:2017ngi ; Guo:2022mpr , stochastic GWs Brito:2017wnc ; Brito:2017zvb and GW beats Guo:2022mpr . The consequence of the BH superradiance could also be indirectly studied via the mass-spin distribution of numerous BHs, which would align along Regge trajectories if superradiance occurs Arvanitaki:2010sy ; Cardoso:2018tly ; Fernandez:2019qbj ; Ng:2019jsx ; Ng:2020ruv ; Cheng:2022jsw .
@@
Both indirect and direct searches require a thorough understanding of the evolution process of the BH-condensate system. In previous studies, accretion is often overlooked for simplicity. The obtained results are then qualified only for isolated BHs, without gas and stars in the neighborhood. Nonetheless, accretion is important in the evolution of most BHs, even crucial for those BHs in active galactic nuclei or X-ray binaries. To our best knowledge, the impact of gas accretion on the superradiant evolution was first pointed out in Ref. Brito:2014wla and further studied in Refs. Fukuda:2019ewf ; Hui:2022sri ; Unal:2023yxt .
@@
In this work, we carefully investigate the effect of accretion on the evolution of BH-condensate systems. We numerically solve the evolution equations with both the dominant and the subdominant modes. Analytical approximations of important quantities and time scales are also provided and compared with the numerical results. We then discuss the effects of accretion on the GW signals, including GW beat, which was proposed to distinguish the GW from superradiance and other monochromatic sources Guo:2022mpr . These numerical and analytical results provide baselines for more complicated cases, such as those with nonzero couplings between the ultralight scalar and the SM particles.
@@
This paper is organized as follows. Sec. II begins with a brief review of scalar superradiance, followed by discussions on the evolution of BHs without the scalar field or accretion. Then in Sec. III, we study the effect of accretion on evolution with solely the dominant mode and provide analytic estimates of the essential quantities. This approach is then generalized to the scenario with multiple modes in Sec. IV, with a brief discussion of the impact of accretion on GW signals. Finally, we summarize our findings in Sec. V.
[…]
II Setup
II.1 Scalar superradiance
@@
The Kerr spacetime metric with mass $M$ and angular momentum $J$ is expressed in the Boyer-Lindquist coordinates as Boyer:1966qh
$ds^{2}=$ $\left(1-\frac{2Mr}{\Sigma}\right)dt^{2}+\frac{4aMr}{\Sigma}\sin^{2}\theta dtd\varphi-\frac{\Sigma}{\Delta}dr^{2}$ (1)
@@
$a$ $\equiv J/M,$ (2a)
$\Delta$ $\equiv r^{2}-2Mr+a^{2},$ (2b)
$\Sigma$ $\equiv r^{2}+a^{2}\cos^{2}\theta.$ (2c)
@@
In this work, we consider a real scalar field surrounding a Kerr BH. Due to the small energy density of the condensate, its backreaction to the metric and self-interaction of the scalar field can be safely ignored Brito:2014wla . They can be included systematically as perturbations if high precision is required, which is beyond the scope of this work. Below we study a free real scalar field on the Kerr metric
$\left(\nabla^{\rho}\nabla_{\rho}+\mu^{2}\right)\Phi=0,$ (4)
where $\mu$ is the scalar mass. Generally, the energy eigenvalues are complex numbers depending on three indices, which are the overtone number $n$, azimuthal number $l$, and magnetic number $m$. The principal number $\bar{n}=n+l+1$ is also widely used instead of $n$. For eigenstate with indices $\{n,l,m\}$, we use $\omega_{nlm}$ and $\Gamma_{nlm}$ to denote the real and imaginary parts of the eigenvalue, respectively. For bounded scalar fields, the $\omega_{nlm}$ can be expressed as a series of $\alpha=M\mu$ Baumann:2019eav
$\omega_{nlm}$ $=\mu\Big{(}1-\frac{\alpha^{2}}{2\bar{n}^{2}}-\frac{\alpha^{4}}{8\bar{n}^{2}}+\frac{f_{\bar{n}l}}{\bar{n}^{3}}\alpha^{4}+\frac{h_{l}a_{*}m}{\bar{n}^{3}}\alpha^{5}+\cdots\Big{)},$ (5)
@@
$f_{\bar{n}l}$ $\equiv-\frac{6}{2l+1}+\frac{2}{\bar{n}},$ (6)
$h_{l}$ $\equiv\frac{16}{2l(2l+1)(2l+2)}.$ (7)
The imaginary part $\Gamma_{nlm}$, which is also called the superradiance rate, has been calculated at the leading order in Ref. Detweiler:1980uk and has been recently improved to next-to-leading order (NLO) in Ref. Bao:2022hew . Both the expressions have compact forms while the latter is much more accurate compared with the numerical calculation in Ref. Dolan:2007mj . In this work, we use the NLO expression, which gives [^1] This expression is rewritten from the result in Ref. Bao:2022hew .
$\begin{split}&\Gamma_{nlm}=\\ &\hskip 18.49411pt-\omega_{1}\left(4\kappa\sqrt{M^{2}-a^{2}}\right)^{2l^{\prime}+1}\frac{\Gamma(n+2l^{\prime}+2)}{n!}\frac{\sinh(2\pi p)}{2\pi}\\ &\times\frac{\left|\Gamma\left(l^{\prime}+1-ip+\sqrt{q-p^{2}}\right)\Gamma\left(l^{\prime}+1+ip+\sqrt{q-p^{2}}\right)\right|^{2}}{\left[\Gamma(2l^{\prime}+1)\Gamma(2l^{\prime}+2)\right]^{2}},\end{split}$ (8)
@@
$\epsilon$ $\equiv-\frac{8\alpha^{2}}{2l+1},$ (9a)
$\begin{split}q&\equiv\frac{8Mr_{+}\omega_{0}(r_{+}\omega_{0}-mM\Omega_{\text{H}})}{r_{+}-r_{-}}\\ &\hskip 56.9055pt-\mu^{2}(r_{+}^{2}+a^{2})+4M^{2}(\mu^{2}-3\omega^{2}_{0}),\end{split}$ (9b)
$\omega_{0}$ $\equiv\mu\sqrt{1-\frac{2\alpha^{2}}{\bar{n}^{2}+4\alpha^{2}+\bar{n}\sqrt{\bar{n}^{2}+8\alpha^{2}}}},$ (9c)
$\omega_{1}$ $\equiv\frac{\mu^{2}-\omega_{0}^{2}}{\bar{n}\omega_{0}(1+4M^{2}(\omega_{0}^{2}-\mu^{2})/\bar{n})}.$ (9d)
Eq. (8) indicates that the superradiance condition $\Gamma_{nlm}>0$ is equivalent to $\omega<m\Omega_{\text{H}}$, from which one could derive the critical BH spin for each mode labeled by $\{n,l,m\}$
$a^{(nlm)}_{\mathrm{*c}}=\frac{4mM\omega_{nlm}}{m^{2}+\left(2M\omega_{nlm}\right)^{2}}\text{, for $M\omega_{nlm}\leq\frac{m}{2}$}.$ (10)
@@
The superradiance rate could be very different for different $\{n,l,m\}$ modes. In general, the most important modes are those with $l=m$ which is the smallest integer greater than $\omega_{nlm}/\Omega_{H}$. Among these modes, the superradiance rate is usually smaller with a larger value of $n$. This hierarchy is strict for $l=1$ and 2. For modes with $l\geq 3$, mode with larger $n$ may have a larger superradiance rate at some ranges of $M\mu$ and $a_{*}$, which is named “overtone mixing” in the literature Siemonsen:2019ebd . In this work, we first focus on the case in which modes with larger $n$ have smaller superradiance rates, obtaining analytical approximations of all important quantities. Then in the last section, we argue that these approximations also work in the presence of “overtone mixing”.
II.2 Evolution equations
@@
The evolution of the BH-condensate system includes the exchange of energy and angular momentum between the BH and the scalar field. The latter emits GW while rotating around the host BH. With accretion, the BH also absorbs energy from the accretion disk. In a more realistic scenario, the photons radiated by the accretion disk are important when the BH spin is very close to identity Thorne:1974ve . The evolution equations with all these effects are
$\dot{M}$ $=\dot{M}_{\mathrm{acc}}\left(1+\frac{E_{\mathrm{rad}}^{\dagger}}{E_{\mathrm{ms}}^{\dagger}}\right)-\sum_{nlm}\dot{E}_{\mathrm{SR}}^{(nlm)},$ (11a)
$\dot{J}$ $=\frac{\dot{M}_{\mathrm{acc}}}{E_{\mathrm{ms}}^{\dagger}}\left(J_{\mathrm{ms}}^{\dagger}+J_{\mathrm{rad}}^{\dagger}\right)-\sum_{nlm}\frac{m}{\omega_{nlm}}\dot{E}_{\mathrm{SR}}^{(nlm)},$ (11b)
$\dot{M}_{\mathrm{s}}^{(nlm)}$ $=\dot{E}_{\mathrm{SR}}^{(nlm)}-\dot{E}_{\mathrm{GW}}^{(nlm)},$ (11c)
$\dot{J}_{\mathrm{s}}^{(nlm)}$ $=\frac{m}{\omega_{nlm}}\left(\dot{E}_{\mathrm{SR}}^{(nlm)}-\dot{E}_{\mathrm{GW}}^{(nlm)}\right),$ (11d)
@@
where $M_{s}^{(nlm)}$, $J_{s}^{(nlm)}$, $\dot{E}_{\mathrm{SR}}^{(nlm)}$ and $\dot{E}_{\mathrm{GW}}^{(nlm)}$ are the mass, angular momentum, superradiance rate and the GW emission rate of the $\{n,l,m\}$ mode, respectively. The $\dot{E}_{\mathrm{SR}}^{(nlm)}$ is related to the imaginary part of the eigenfrequency by $\dot{E}_{\mathrm{SR}}^{(nlm)}\equiv 2{M}_{\mathrm{s}}^{(nlm)}\Gamma_{nlm}$. The $\dot{M}_{\mathrm{acc}}$ is the mass absorption rate of the BH from the accretion disk. The $E_{\mathrm{ms}}^{\dagger}$ and $J_{\mathrm{ms}}^{\dagger}$ are the energy and angular momentum of unit mass in the last stable circular orbit, respectively. The accretion disk radiates photons, some of which are later captured by the BH. Its contribution to the BH mass and angular momentum are described by $E_{\mathrm{rad}}^{\dagger}$ and $J_{\mathrm{rad}}^{\dagger}$ for unit accreted mass, respectively, which are calculated in Refs. Page:1974he ; Thorne:1974ve . The expressions for $E_{\mathrm{ms}}^{\dagger}$, $J_{\mathrm{ms}}^{\dagger}$ and $\dot{E}_{\mathrm{GW}}^{(nlm)}$ are listed in App. A for convenience.
In obtaining Eqs. (11), we have made the following assumptions:
@@
  • The backreactions of the condensate and the accretion disk on the metric, as well as the scalar self-interaction are ignored due to the low energy densities of the condensate and the disk Brito:2014wla .
@@
  • The quasi-adiabatic approximation is employed, which assumes that the dynamical timescale of the BH is much shorter than the timescales of both the accretion and superradiant instability Brito:2014wla ; Brito:2017zvb .
@@
II.2.1 The evolution without the scalar field
In this part, we study the scenario in the absence of the scalar field. If the photon radiation can be further ignored, i.e. $E_{\mathrm{rad}}^{\dagger}=J_{\mathrm{rad}}^{\dagger}=0$, the evolution of the BH mass and angular momentum only depends on the accretion rate $\dot{M}_{\text{acc}}$. Useful results can be obtained even without any knowledge of $\dot{M}_{\text{acc}}$. From Eqs. (11a) and (11b), one could derive
$\frac{\mathrm{d}a_{*}}{\mathrm{d}\ln M}$ $=\frac{J_{\mathrm{ms}}^{\dagger}}{ME_{\mathrm{ms}}^{\dagger}}-2a_{*}.$ (12)
BHs with constant spin $a_{*}=1$ and any value of $M$ is a trivial solution. For BH with initial spin $a_{*0}<1$ and mass $M_{0}$, a non-trivial solution also exist Thorne:1974ve
$a_{*}=\begin{cases}\sqrt{\frac{2}{3}}\frac{kM_{0}}{M}\left(4-\sqrt{\frac{18k^{2}M_{0}^{2}}{M^{2}}-2}\right),&\text{if\ }1\leq\frac{M}{M_{0}}\leq{\sqrt{6}\,k},\\ 1,&\text{if\ }\frac{M}{M_{0}}\geq{\sqrt{6}\,k},\end{cases}$ (13)

### v2

# Effect of accretion on scalar superradiant instability

@@ 1 Introduction
Superradiance allows ultra-light bosons to extract energy and angular momentum from rotating black holes (BHs). This mechanism exists when the condition $\omega<m\Omega_{\mathrm{H}}$ is satisfied, where $\omega$ is the frequency of the boson field, $m$ is the magnetic number, and $\Omega_{\mathrm{H}}$ is the angular velocity of the event horizon. This process leads to a significant accumulation of bosons in bound states surrounding a Kerr BH, forming a BH-condensate system analogous to a hydrogen atom. For a comprehensive review on superradiance, we refer the readers to Ref. [1]. In this work, we study the superradiance of ultra-light scalars, such as the QCD axion [2, 3], and axion-like particles suggested by string theory [4].
@@
Many observables are proposed to observe such BH-condensate systems [5, 6, 7]. In this article, we focus on the gravitational effects of the scalars, independent of how the scalar field couples to the Standard Model particles. Both direct and indirect detection methods have been proposed in the literature. One could directly observe the gravitational wave (GW) signals from the BH-condensate system, including quasi-monochromatic GWs [8, 9, 10, 11, 12, 13, 14, 15, 16], stochastic GWs [13, 14] and GW beats [16]. The consequence of the BH superradiance could also be indirectly studied via the mass-spin distribution of numerous BHs, which would align along Regge trajectories if superradiance occurs [8, 17, 18, 19, 20, 21].
@@
Both indirect and direct searches require a thorough understanding of the evolution process of the BH-condensate system. In previous studies, accretion is often overlooked for simplicity. The obtained results are then qualified only for isolated BHs, without gas and stars in the neighborhood. Nonetheless, accretion is important in the evolution of most BHs, even crucial for those BHs in active galactic nuclei or X-ray binaries. To our best knowledge, the impact of gas accretion on the superradiant evolution was first pointed out in Ref. [22] and further studied in Refs. [23, 24, 25, 26, 27].
@@
In this work, we carefully investigate the effect of accretion on the evolution of BH-condensate systems. We numerically solve the evolution equations with both the dominant and the subdominant modes. Analytical approximations of important quantities and time scales are also provided and compared with the numerical results. We then discuss the effects of accretion on the GW signals, including GW beat, which was proposed to distinguish the GW from superradiance and other monochromatic sources [16]. These numerical and analytical results provide baselines for more complicated cases, such as those with nonzero couplings between the ultralight scalar and the SM particles.
@@
This paper is organized as follows. Sec. 2 begins with a brief review of scalar superradiance, followed by discussions on the evolution of BHs without the scalar field or accretion. Then in Sec. 3, we study the effect of accretion on evolution with solely the dominant mode and provide analytic estimates of the essential quantities. This approach is then generalized to the scenario with multiple modes in Sec. 4, with a brief discussion of the impact of accretion on GW signals. Finally, we summarize our findings in Sec. 5.
Throughout the paper, we adopt the Planck units in which $G=\hbar=c=1$, and use the $\{+,-,-,-\}$ signature.
2 Setup
2.1 Scalar superradiance
@@
The Kerr spacetime metric with mass $M$ and angular momentum $J$ is expressed in the Boyer-Lindquist coordinates as [28]
$ds^{2}=$ $\left(1-\frac{2Mr}{\Sigma}\right)dt^{2}+\frac{4aMr}{\Sigma}\sin^{2}\theta dtd\varphi-\frac{\Sigma}{\Delta}dr^{2}$ (2.1)
@@
$a$ $\equiv J/M,$ (2.2a)
$\Delta$ $\equiv r^{2}-2Mr+a^{2},$ (2.2b)
$\Sigma$ $\equiv r^{2}+a^{2}\cos^{2}\theta.$ (2.2c)
@@
In this work, we consider a real scalar field surrounding a Kerr BH. Due to the small energy density of the condensate, its backreaction to the metric and self-interaction of the scalar field can be safely ignored [22]. They can be included systematically as perturbations if high precision is required, which is beyond the scope of this work. Below we study a free real scalar field on the Kerr metric
$\left(\nabla^{\rho}\nabla_{\rho}+\mu^{2}\right)\Phi=0,$ (2.4)
where $\mu$ is the scalar mass. Generally, the energy eigenvalues are complex numbers depending on three indices, which are the overtone number $n$, azimuthal number $l$, and magnetic number $m$. The principal number $\bar{n}=n+l+1$ is also widely used instead of $n$. For eigenstate with indices $\{n,l,m\}$, we use $\omega_{nlm}$ and $\Gamma_{nlm}$ to denote the real and imaginary parts of the eigenvalue, respectively. For bounded scalar fields, the $\omega_{nlm}$ can be expressed as a series of $\alpha=M\mu$ [29]
$\omega_{nlm}$ $=\mu\Big(1-\frac{\alpha^{2}}{2\bar{n}^{2}}-\frac{\alpha^{4}}{8\bar{n}^{2}}+\frac{f_{\bar{n}l}}{\bar{n}^{3}}\alpha^{4}+\frac{h_{l}a_{*}m}{\bar{n}^{3}}\alpha^{5}+\cdots\Big),$ (2.5)
@@
$f_{\bar{n}l}$ $\equiv-\frac{6}{2l+1}+\frac{2}{\bar{n}},$ (2.6)
$h_{l}$ $\equiv\frac{16}{2l(2l+1)(2l+2)}.$ (2.7)
The imaginary part $\Gamma_{nlm}$, which is also called the superradiance rate, has been calculated at the leading order in Ref. [30] and has been recently improved to next-to-leading order (NLO) in Ref. [31]. Both the expressions have compact forms while the latter is much more accurate compared with the numerical calculation in Ref. [32]. In this work, we use the NLO expression, which gives [^1] This expression is rewritten from the result in Ref. [31].
$\begin{split}&\Gamma_{nlm}=-\omega_{1}\left(4\kappa\sqrt{M^{2}-a^{2}}\right)^{2l^{\prime}+1}\frac{\Gamma(n+2l^{\prime}+2)}{n!}\frac{\sinh(2\pi p)}{2\pi}\\ &\hskip 34.14322pt\times\frac{\left|\Gamma\left(l^{\prime}+1-ip+\sqrt{q-p^{2}}\right)\Gamma\left(l^{\prime}+1+ip+\sqrt{q-p^{2}}\right)\right|^{2}}{\left[\Gamma(2l^{\prime}+1)\Gamma(2l^{\prime}+2)\right]^{2}},\end{split}$ (2.8)
@@
$\epsilon$ $\equiv-\frac{8\alpha^{2}}{2l+1},$ (2.9a)
$q$ $\equiv\frac{8Mr_{+}\omega_{0}(r_{+}\omega_{0}-mM\Omega_{\text{H}})}{r_{+}-r_{-}}-\mu^{2}(r_{+}^{2}+a^{2})+4M^{2}(\mu^{2}-3\omega^{2}_{0}),$ (2.9b)
$\omega_{0}$ $\equiv\mu\sqrt{1-\frac{2\alpha^{2}}{\bar{n}^{2}+4\alpha^{2}+\bar{n}\sqrt{\bar{n}^{2}+8\alpha^{2}}}},$ (2.9c)
$\omega_{1}$ $\equiv\frac{\mu^{2}-\omega_{0}^{2}}{\bar{n}\omega_{0}(1+4M^{2}(\omega_{0}^{2}-\mu^{2})/\bar{n})}.$ (2.9d)
Eq. (2.8) indicates that the superradiance condition $\Gamma_{nlm}>0$ is equivalent to $\omega<m\Omega_{\text{H}}$, from which one could derive the critical BH spin for each mode labeled by $\{n,l,m\}$
$a^{(nlm)}_{\mathrm{*c}}=\frac{4mM\omega_{nlm}}{m^{2}+\left(2M\omega_{nlm}\right)^{2}}\text{, for $M\omega_{nlm}\leq\frac{m}{2}$}.$ (2.10)
@@
The superradiance rate could be very different for different $\{n,l,m\}$ modes. In general, the most important modes are those with $l=m$ which is the smallest integer greater than $\omega_{nlm}/\Omega_{H}$. Among these modes, the superradiance rate is usually smaller with a larger value of $n$. This hierarchy is strict for $l=1$ and 2. For modes with $l\geq 3$, mode with larger $n$ may have a larger superradiance rate at some ranges of $M\mu$ and $a_{*}$, which is named “overtone mixing” in the literature [33]. In this work, we first focus on the case in which modes with larger $n$ have smaller superradiance rates, obtaining analytical approximations of all important quantities. Then in the last section, we argue that these approximations also work in the presence of “overtone mixing”.
2.2 Evolution equations
@@
The evolution of the BH-condensate system includes the exchange of energy and angular momentum between the BH and the scalar field. The latter emits GW while rotating around the host BH. With accretion, the BH also absorbs energy from the accretion disk. In a more realistic scenario, the photons radiated by the accretion disk are important when the BH spin is very close to 1 [34]. The evolution equations with all these effects are [22, 34]
$\dot{M}$ $=\dot{M}_{\mathrm{acc}}\left(1+\frac{E_{\mathrm{rad}}^{\dagger}}{E_{\mathrm{ms}}^{\dagger}}\right)-\sum_{nlm}\dot{E}_{\mathrm{SR}}^{(nlm)},$ (2.11a)
$\dot{J}$ $=\frac{\dot{M}_{\mathrm{acc}}}{E_{\mathrm{ms}}^{\dagger}}\left(J_{\mathrm{ms}}^{\dagger}+J_{\mathrm{rad}}^{\dagger}\right)-\sum_{nlm}\frac{m}{\omega_{nlm}}\dot{E}_{\mathrm{SR}}^{(nlm)},$ (2.11b)
$\dot{M}_{\mathrm{s}}^{(nlm)}$ $=\dot{E}_{\mathrm{SR}}^{(nlm)}-\dot{E}_{\mathrm{GW}}^{(nlm)},$ (2.11c)
$\dot{J}_{\mathrm{s}}^{(nlm)}$ $=\frac{m}{\omega_{nlm}}\left(\dot{E}_{\mathrm{SR}}^{(nlm)}-\dot{E}_{\mathrm{GW}}^{(nlm)}\right),$ (2.11d)
@@
where $M_{s}^{(nlm)}$, $J_{s}^{(nlm)}$, $\dot{E}_{\mathrm{SR}}^{(nlm)}$ and $\dot{E}_{\mathrm{GW}}^{(nlm)}$ are the mass, angular momentum, superradiance rate and the GW emission rate of the $\{n,l,m\}$ mode, respectively. The $\dot{E}_{\mathrm{SR}}^{(nlm)}$ is related to the imaginary part of the eigenfrequency by $\dot{E}_{\mathrm{SR}}^{(nlm)}\equiv 2{M}_{\mathrm{s}}^{(nlm)}\Gamma_{nlm}$. The $\dot{M}_{\mathrm{acc}}$ is the mass absorption rate of the BH from the accretion disk. The $E_{\mathrm{ms}}^{\dagger}$ and $J_{\mathrm{ms}}^{\dagger}$ are the energy and angular momentum of unit mass in the last stable circular orbit, respectively. The accretion disk radiates photons, some of which are later captured by the BH. Its contribution to the BH mass and angular momentum are described by $E_{\mathrm{rad}}^{\dagger}$ and $J_{\mathrm{rad}}^{\dagger}$ for unit accreted mass, respectively, which are calculated in Refs. [35, 34]. The expressions for $E_{\mathrm{ms}}^{\dagger}$, $J_{\mathrm{ms}}^{\dagger}$ and $\dot{E}_{\mathrm{GW}}^{(nlm)}$ are listed in App. A for convenience.
In obtaining Eqs. (2.11), we have made the following assumptions:
@@
  • The backreactions of the condensate and the accretion disk on the metric, as well as the scalar self-interaction are ignored due to the low energy densities of the condensate and the disk [22].
@@
  • The quasi-adiabatic approximation is employed, which assumes that the dynamical timescale of the BH is much shorter than the timescales of both the accretion and superradiant instability [22, 14].
@@
2.2.1 The evolution without the scalar field
In this part, we study the scenario in the absence of the scalar field. If the photon radiation can be further ignored, i.e. $E_{\mathrm{rad}}^{\dagger}=J_{\mathrm{rad}}^{\dagger}=0$, the evolution of the BH mass and angular momentum only depends on the accretion rate $\dot{M}_{\text{acc}}$. Useful results can be obtained even without any knowledge of $\dot{M}_{\text{acc}}$. From Eqs. (2.11a) and (2.11b), one could derive
$\frac{\mathrm{d}a_{*}}{\mathrm{d}\ln M}$ $=\frac{J_{\mathrm{ms}}^{\dagger}}{ME_{\mathrm{ms}}^{\dagger}}-2a_{*}.$ (2.12)
BHs with constant spin $a_{*}=1$ and any value of $M$ is a trivial solution. For BH with initial spin $a_{*0}<1$ and mass $M_{0}$, a non-trivial solution also exist [34]
$a_{*}=\begin{cases}\sqrt{\frac{2}{3}}\frac{kM_{0}}{M}\left(4-\sqrt{\frac{18k^{2}M_{0}^{2}}{M^{2}}-2}\right),&\text{if\ }1\leq\frac{M}{M_{0}}\leq{\sqrt{6}\,k},\\ 1,&\text{if\ }\frac{M}{M_{0}}\geq{\sqrt{6}\,k},\end{cases}$ (2.13)